# expected utility

### v1

@@ 6 Conclusion
In this paper, we derived explicit analytic formulas for general contracts, including participating life insurance contracts, when optimizing mean-variance in the multi-dimensional Black-Scholes market. Moreover, by showing the existence of all arising parameters, we showed the existence of the optimal solution. We also gave the HJB equation for the value functional if the market is possibly incomplete. A numerical analysis shows that the mean-variance optimal strategy compared to the EU optimal strategy becomes more conservative the shorter the maturity is and is increased in particular in bad economic states. Future research directions include possibly generalizing this approach to other stock market models.

### v2

@@ 5 Conclusion
In this paper, we derived explicit analytic formulas for general contracts, including participating life insurance contracts, when optimizing mean-variance in the multi-dimensional Black-Scholes market. Moreover, by showing the existence of all arising parameters, we showed the existence of the optimal solution. A numerical analysis shows that the mean-variance optimal strategy compared to the EU optimal strategy becomes more conservative the shorter the maturity is and is increased in particular in bad economic states. Future research directions include possibly generalizing this approach to other stock market models.
@@ Appendix A Proofs
  • Proof of Proposition 3.6.
    For this proof, we show that the alternative formulation of this proposition admits the same Lagrange function as our main problem given in (3.5). Therefore, we first reformulate the problem as a maximization problem by replacing $\min_{u\in\mathcal{U}}\mathbb{E}[\tilde{L}(X_{T})]$ with $\max_{u\in\mathcal{U}}\mathbb{E}[-\tilde{L}(X_{T})]$. Then, we get for the Lagrange function with $X\geq 0$:
    $L(X,y,\lambda)$ $=-\tilde{L}(X)-\lambda[\alpha k_{0}-\alpha(X-k_{1})\mathbbm{1}_{k_{1}\leq X<k_{2}}-\tilde{\alpha}(X-k_{2}+\tfrac{\alpha}{\tilde{\alpha}}(k_{2}-k_{1}))\mathbbm{1}_{X\geq k_{2}}]-y\xi X$
    =$=\begin{cases}-\gamma\alpha^{2}k_{0}^{2}-\lambda\alpha k_{0}-y\xi X,&\text{if }X\in[0,k_{1}),\\ -\gamma\alpha^{2}(X-k_{1}-k_{0})^{2}-\lambda\alpha k_{0}+\lambda\alpha(X-k_{1})-y\xi X,&\text{if }X\in[k_{1},k_{2}),\\ -\gamma\tilde{\alpha}^{2}(X-k_{2}+\tfrac{\alpha}{\tilde{\alpha}}(k_{2}-k_{1}-k_{0}))^{2}-\lambda\alpha k_{0}&\\ \hskip 85.35826pt+\lambda\tilde{\alpha}(X-k_{2}+\tfrac{\alpha}{\tilde{\alpha}}(k_{2}-k_{1}))-y\xi X,&\text{if }X\in[k_{2},\infty),\end{cases}$
    =$=\begin{cases}-\gamma\alpha^{2}k_{0}^{2}-\lambda\alpha k_{0}-y\xi X,&\text{if }X\in[0,k_{1}),\\ -\gamma\alpha^{2}(X-k_{1}-k_{0})^{2}+\lambda\alpha(X-k_{1}-k_{0})-y\xi X,&\text{if }X\in[k_{1},k_{2}),\\ -\gamma(\tilde{\alpha}(X-k_{2})+\alpha(k_{2}-k_{1}-k_{0}))^{2}&\\ \hskip 85.35826pt+\lambda(\tilde{\alpha}(X-k_{2})+\alpha(k_{2}-k_{1}-k_{0}))-y\xi X,&\text{if }X\in[k_{2},\infty).\end{cases}$
    By comparing the terms with (3.5), we observe that the Lagrange functions are identical which gives us the proof. ∎